A charged particle of mass m and change q is released form rest in an electric field of constant magnitude e .the kinetic energy of the particle after time t is

Grade:12

1 Answers

Sounak Dutta
6 years ago
According to the question,
Force= q*e
so acceleration is (q*e)/m
after time t
velocity is (q*e*t)/m
so kinetic energy is 1/2*m*((q*e*t)/m)^2=1/2*((q*e*t)^2)/m
Thanks